A wheel rotates with a constant angular acceleration of 3.5 rad/s2. If the angular speed of the wheel is 2 rad/s at t = 0, Find what angle does the wheel rotate between t = 0 and t = 2 s ?




1
Expert's answer
2022-03-24T11:00:50-0400

The total angle the wheel rotate between t = 0 and t = 2 s

"\theta=\omega_0t+\beta t^2/2\\
=2*2+3.5*2^2/2=11\:\rm rad"